OpenStudy (anonymous):

Determine quadratic function given vertex and a point: Vertex: (1,1) and point: (2,4) Please explain Thanks!

OpenStudy (anonymous):

Go back to the form: \[y=ax ^{2}+bx+c\] Now, the x-coord of the vertex = -b/2a, so \[1=\frac{ -b }{ 2a }\] Then: \[\frac{ 1 }{ 1 }=\frac{ -b }{ 2a }\] cross multiply and we get b = -2a

OpenStudy (anonymous):

Now we can replace b with 2a: \[y=ax ^{2}+2ax+c\] Plug in the first point replacing x with 1 and y with 1 and you get an equation with just 2 unknowns: a and c Plug in the second point replacing the x and y again and you get a second equation with just 2 unknowns: a and c Now you can solve this system of equations as you have 2 unknowns and 2 equations.

OpenStudy (anonymous):

Do you know how to solve a system of equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry, i didn't get a message when you answered Did you get: 3a + c = 1 8a + c = 4 ?

OpenStudy (anonymous):

To solve a system, we want to be able to add 2 equations and have one of the variables "disappear". We could multiply that first equation by (-1). We get: -3a - c = -1 8a + c = 4 Notice that the signs in the first equation all changed because we mult by -1 Now add the equations downward. The c's should disappear.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I made an error.\[-\frac{ b }{ 2a } = 1\] So b = -2a ... sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[y=ax ^{2}-2ax+c\]

OpenStudy (anonymous):

when we plug in (1,1), we get: -a + c = 1 When we plug in (2, 4), we get: 4 = 4a - 4a + c So: 4 = c Now you can use that to find a since: -a + c = 1 and c = 4, -a + 4 = 1 so a = ?

OpenStudy (anonymous):

also, b = -2a, and now that we know 'a' we can use it to find 'b'.

OpenStudy (anonymous):

would it be -2a?

OpenStudy (anonymous):

-a + 4 = 1 a=3

OpenStudy (anonymous):

so now what did you get for 'a' and 'b'?

OpenStudy (anonymous):

sorry, 'c' and 'b'

OpenStudy (anonymous):

a=3,b=-2a

OpenStudy (anonymous):

c=4, b=2a?

OpenStudy (anonymous):

b = -2a = -2(3) = _?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

remember I made that error and b = - 2a, not 2a

OpenStudy (anonymous):

Good then you go back to \[y=ax ^{2}+bx+c \] and put 3 in for 'a', -6 in for b and 4 in for c. That gives you this specific quadratic equation A lot of work on that one!

OpenStudy (anonymous):

I know thank you so much I need to work on my math skills more

OpenStudy (anonymous):

if you plug in the given points to your resulting equation, you will find they both work. If you find the x-coordinate of the vertex, you will find it is 1

OpenStudy (anonymous):

good luck I am a math teacher / tutor at a community college but i have developed nerve issues so I had to cut back and this is fun for me

OpenStudy (anonymous):

What college is that? I should attend